Bailey, Inc., is considering buying a new gang punch that would allow them to produce circuit boards more efficiently. The punch has a first cost of $60,000 and a useful life of 15 years. At the end of its useful life, the punch has no salvage value. Annual labor costs would increase $3,500 using the gang punch, but annual raw material costs would decrease $7,000. MARR is 4.75 %/year.

1.What is the present worth of this investment?

2.What is the decision rule for judging the attractiveness of investments based on present worth?

3.Should Bailey buy the gang punch?
Answer #1

initial cost = 60000

MARR = 4.75%

t = 15 yrs

Annual savings = 7000

Annual cost = 3500

Net annual revenue = 7000 - 3500 = 3500

PW = -60000 + 3500*(P/A, 4.75%,15)

= -60000 + 3500 * 10.55730

= -23049.45

Decision rule : when PW>0, project should be selected

No as PW<0